OpenStudy (anonymous):

Solve 6 over x minus 3 equals 3 over x for x and determine if the solution is extraneous or not.

OpenStudy (anonymous):

x is 1

OpenStudy (anonymous):

I got x=-3

OpenStudy (anonymous):

6/x - 3 = 3/x 6/x -3/x = 3 3/x = 3 3/3 = x 1= x

OpenStudy (anonymous):

ooooooo ok thanks